OpenStudy (anonymous):

Solve the equation 3y(y+4)=15

OpenStudy (radar):

Divide thru by 3 getting: y(y+4)=5 y^2+4y-5=0 factor (y+5)(y-1)= 0 solve y=1 y=-5

OpenStudy (anonymous):

first of all multiply the 3y to the brackets so it will be 3y*y=3y^2 ,3y*4=12y and then enter the 15 in the other side : it will be 3y^2+12y-15 and then u can take 3 common : 3 (y^2+4y-5) and then facorize : 3 (y-1) (y+5)